# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v2

@@ 2 Special elements in the image of the convolution of two modular forms
For (vi), let $r\geq 1$ be such that $\varepsilon_{f}^{2^{r}}$ is the trivial character. Let $(\gamma,\chi)$ be such that $f$ has an inner twist by $(\gamma,\chi)$. Then, by Lemma 1, $\chi$ is the product of a quadratic character and a power of $\varepsilon_{f}$, so $\chi^{2^{r}}$ is trivial, so $\mathrm{Gal}(H/{\mathbb{Q}})$ has exponent dividing $2^{r}$. Therefore, $d:=[H:{\mathbb{Q}}]$ is a power of $2$.
We want to apply Proposition 2, but we need to rule out its exceptional cases. The first one cannot occur since, under our assumptions, $[H:{\mathbb{Q}}]$ is not divisible by $3$.
Let $I$ denote the image in $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ of $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})$. The extension $H/{\mathbb{Q}}$ is Galois of degree $d$, so one has $I\triangleleft\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. Since $d$ kills the quotient $\rho_{g}(G_{{\mathbb{Q}}})/\rho_{g}(G_{H})$, $d$ kills the subgroup $A_{4}/I$. Since $A_{4}$ has exponent $6$ and $d$ is a power of $2$, the group $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/I$ is killed by $2$, hence it is abelian. The abelialization of $A_{4}$ is cyclic of order $3$, so $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/I$ is trivial and $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is surjective, which rules out the second special case of Proposition 2. $\Box$
@@
Remark  When checking the Bloch–Kato conjecture, it seems more appropriate to use the $L$-function $L^{aut}(E,\rho,s)$ attached to the Rankin–Selberg convolution of the unitary automorphic representations attached to $f$ and $g$ (as in [2, Section 3.6]), since it admits a holomorphic continuation and satisfies a functional equation (with center $\frac{1}{2}$, see [6, Theorem 11.7.1]). We are using the imprimitive $L$-function in the statement above primarily for the sake of simplicity, since it is the one appearing throughout [10]. In fact, using the classical description of $L^{aut}(E,\rho,s)$ given in [11, Theorem 2.2], it is not difficult to check that the meromorphic continuation of $L(E,\rho,s)$ is indeed holomorphic near $s=1$ and that one has $L(E,\rho,1)\neq 0$ if and only if $L^{aut}(E,\rho,\frac{1}{2})\neq 0$.
@@ 3 Families of dihedral counter-examples: proof of Theorem Introduction
We first recall the following fact, already identified by Loeffler as an obstruction in [14, Proposition 4.1.1].
@@
Thus, we can (and do) assume that $M$ and $N$ both have distinct eigenvalues, respectively $m,m^{\prime}$ and $n,n^{\prime}$. Then $M\otimes N$ is diagonalizable with eigenvalues $mn,m^{\prime}n^{\prime},m^{\prime}n,nm^{\prime}$. Since $1=(mn)(m^{\prime}n^{\prime})=(mn^{\prime})(m^{\prime}n)$, there is never exactly one of these eigenvalues which is equal to one. $\Box$
As noted by Loeffler, this lemma implies that, for any couple $(f^{\prime},g^{\prime})$ of newforms, $T_{f^{\prime},g^{\prime},\mathfrak{p}}$ cannot satisfy (wE) if the product of the characters of $f^{\prime},g^{\prime}$ is trivial, whence the requirement that $\varepsilon_{f}\varepsilon_{g}\neq 1$ in Question Introduction.
@@
    starts3-2special Let $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ be a prime ideal and $\sigma\in G_{{\mathbb{Q}}}$ be such that the kernel of $\sigma-1$ acting on $T_{f,g,\mathfrak{p}}$ is a line. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}})$ and $N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be matrices for the action of $\sigma$ on $T_{f,\mathfrak{p}},T_{g}$. Then one has $\varepsilon(\sigma)=\varepsilon_{f}(\sigma)\varepsilon_{g}(\sigma)=\det{M}\det{N}\neq 1$ by Lemma 3, so $\sigma\notin G_{K}$. Since $g$ has complex multiplication by $K$, $N$ has trace zero, so, if $u\in{\mathbb{C}}^{\times}$ denotes a square root of $-\det{N}$, $N$ is similar to $\Delta_{u,-u}$. This implies that $N^{2}$ is a scalar matrix. Moreover, $N$ is similar to $-N$, so $M\otimes N$ is similar to $M\otimes(-N)=-(M\otimes N)$, so that $\ker(M\otimes N+1)$ is also a line.
@@
Proof. –  Suppose that $T_{f,g,\mathfrak{p}}$ satisfies (wE). Let $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ be such that $\ker(\sigma-1\mid T_{f,g,\mathfrak{p}})$ is a line. Let $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be associated to $\sigma$ by Corollary 2. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}}),N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be the matrices given by the action of $\sigma$ on $T_{f,\mathfrak{p}}$ and $T_{g}$, respectively. By point (ii) in the Remarks following Lemma 3, one has $\sigma\notin G_{K}$ and the matrix $N^{2}$ is scalar. In fact, one has by Cayley–Hamilton $N^{2}=(-\det{N})I_{2}=(-\varepsilon_{g}(\sigma))I_{2}=(-\varepsilon(\sigma)\varepsilon_{f}(\sigma)^{-1})I_{2}=\varepsilon_{f}(\sigma)^{-1}I_{2}$. Let $\omega\in L_{0}$ be a square root of $\varepsilon_{f}(\sigma)^{-1}$, then the two eigenvalues of $N$ are $\omega$ and $-\omega$.
@@
By Corollary 2, we can write $M=\alpha^{-1}\varepsilon_{f}(\sigma)M^{\prime}$, where $M^{\prime}\in\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ has determinant $\frac{\alpha^{2}}{\varepsilon_{f}(\sigma)}$. Thus $\alpha\omega^{2}$ is an eigenvalue of $M^{\prime}\otimes\Delta_{\omega,-\omega}$, so that one of the eigenvalues of $M^{\prime}$ is ±$\pm\omega\alpha$. Since (±$(\pm\omega\alpha)^{2}=\det{M^{\prime}}$, one has ±$\pm\alpha\omega=\frac{1}{2}\operatorname{Tr}{M^{\prime}}\in F_{\mathfrak{p}}$, so $\mathcal{B}(\sigma,\mathrm{Frob}_{\mathfrak{p}})=1\neq\varepsilon(\sigma)$.
@@
Suppose conversely that $\mathcal{B}(\cdot,\mathrm{Frob}_{\mathfrak{p}})\neq\varepsilon$. Since $\mathcal{B}_{\mathfrak{p}}^{2}=1$ and $\varepsilon\neq 1$, $\varepsilon$ is not a power of $\mathcal{B}_{\mathfrak{p}}$, so $\ker{\mathcal{B}(\cdot,\mathrm{Frob}_{\mathfrak{p}})}$ is not contained in $\ker{\varepsilon}$. Hence, there exists $s\in\mathrm{Gal}(H/{\mathbb{Q}})$ such that $\varepsilon(s)=-1$ and $\mathcal{B}_{\mathfrak{p}}(s)=1$. By Proposition 2 let $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ be such that its image in $\mathrm{Gal}(H/{\mathbb{Q}})$ is $s$. Let $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be attached to $\sigma$ by Corollary 2 and $\nu:=\frac{\alpha^{2}}{\varepsilon_{f}(\sigma)}\in\mathcal{O}_{F,\mathfrak{p}}^{\times}$. Since $\mathcal{B}_{\mathfrak{p}}(s)=1$, $\nu$ is a square in $F_{\mathfrak{p}}$. Therefore, $\varepsilon_{f}(\sigma)^{-1}=\nu\alpha^{-2}$ has a square root $\omega$ in $L_{\mathfrak{p}}$.
@@
Still by Corollary 2, we can choose $\sigma_{1}\in\sigma G_{H(\mu_{p^{\infty}})}$ such that $\sigma_{1}$ acts on $T_{f,\mathfrak{p}}$ by the matrix $\alpha^{-1}\varepsilon_{f}(\sigma)\begin{pmatrix}\nu&1\\ 0&\nu\end{pmatrix}$. Because $\varepsilon(\sigma_{1})=\varepsilon(\sigma)=-1$, the matrix $\tilde{\rho}_{g}(\sigma_{1})\in\operatorname{PGL}_{2}(\mathcal{O}_{L_{0}})$ has order two. Since $\det{\rho_{g}(\sigma_{1})}=-\varepsilon_{f}(\sigma_{1})^{-1}=-\varepsilon_{f}(\sigma)^{-1}$ and $p\neq 2$, $\rho_{g}(\sigma_{1})$ is $\operatorname{GL}_{2}(\mathcal{O}_{(L_{0}){\mathfrak{p}}})$-conjugate to $\Delta_{\omega,-\omega}$. Therefore, $\sigma_{1}$ acts on $T_{f,g,\mathfrak{p}}$ by $\begin{pmatrix}1&a&0&0\\ 0&1&0&0\\ 0&0&-1&\ast\\ 0&0&0&-1\end{pmatrix}$ with $a\in\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\times}$: hence $T_{f,g,\mathfrak{p}}/(\sigma_{1}-1)T_{f,g,\mathfrak{p}}$ is free of $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-rank one. $\Box$
@@
Proof. –  By Chebotarev and the previous Proposition, the answer to Question Introduction for $(f,g)$ is positive if and only if $\varepsilon$ cannot be written as a $\mathcal{B}(\cdot,\tau)$ for any $\tau\in G_{F}$. Because $\mathcal{B}$ is bilinear with values in a ${\mathbb{F}}_{2}$-vector space, the set of the $\mathcal{B}(\cdot,\tau)$ for $\tau\in G_{F}$ is exactly the set of homomorphisms $f:\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\{\pm 1\}$ such that $f(M)=1$, whence the conclusion. $\Box$
@@ 1. $(a)$
    $k$ is even and $H$ is not totally real,
@@ 3. $(c)$
    weps-cspecial $k$ is odd, $[H:H_{0}]=2$, and $\varepsilon_{f}$ does not take its values in $L^{\times 2}$ (this last condition is satisfied if $L$ does not contain a fourth root of unity $i$),
@@ 4. $(d)$
    weps-dspecial $k$ is odd, $[H:H_{0}]=2$, $i\in L$, and for all $(\gamma,\delta)\in\Gamma_{f}$, $\gamma(i)=\delta(-1)i$,
@@ 5. $(e)$
Moreover, when $k$ is odd, $[H:H_{0}]=2$, $i\in L$ but is not a value of $\varepsilon_{f}$, then the existence of such a $\varepsilon$ is equivalent to the last clause of $(e)$. Under the same assumptions, if $\varepsilon_{f}^{2}=1$, then $(d)$ and $(e)$ are equivalent.
Proof. –  Remember that $L$ is the number field generated by the coefficients of $f$, so it is a subfield of ${\mathbb{C}}$. Note that $\mathcal{B}$ realizes an isomorphism from $\mathrm{Gal}(H/{\mathbb{Q}})/M$ to a subgroup $B$ of the ${\mathbb{F}}_{2}$-vector space $\mathrm{Hom}(G_{F},{\mathbb{F}}_{2})$, so $\mathrm{Gal}(H/{\mathbb{Q}})/M$ has exponent $2$.
Suppose first that $k$ is even and $H$ is not totally real. Let $c\in\mathrm{Gal}(H/{\mathbb{Q}})$ be the image of the complex conjugation and let $\alpha\in L^{\times}$ be associated to $c$ as in Corollary 2. It is enough to prove that $c\notin M$. Since $k$ is even, $\varepsilon_{f}(c)=1$, so one has $\mathcal{B}(c,\tau)=\frac{\tau(\alpha)}{\alpha}$ for all $\tau\in G_{F}$. In particular, for every $(\gamma,\delta)\in\Gamma_{f}$, one has $\mathcal{B}(c,\gamma)=\delta(-1)$. Since $H$ is not totally real, there is some $\delta$ such that $\delta(-1)=-1$, so $c\notin M$, whence the conclusion.
Now suppose that $k$ is odd. The goal is to show that there exists a non-trivial character $\mathrm{Gal}(H/{\mathbb{Q}})/\langle M,c\rangle\rightarrow\{\pm 1\}$. To do that, we need to show that $\mathrm{Gal}(H/{\mathbb{Q}})\supsetneq\langle M,c\rangle$, or in other words that $B\supsetneq\{1,\mathcal{B}(c,\cdot)\}$.
Let $\sigma\in M$ be such that $\varepsilon_{f}(\sigma)=1$: let $\alpha\in L^{\times}$ be associated to $\sigma$ by Corollary 2 and $\nu=\frac{\alpha^{2}}{\varepsilon_{f}(\sigma)}=\alpha^{2}\in F^{\times}$. Then, for all $\tau\in G_{F}$, $1=\mathcal{B}(\sigma,\tau)=\frac{\tau(\sqrt{\nu})}{\sqrt{\nu}}=\frac{\tau(\alpha)}{\alpha}$, hence $\alpha\in F^{\times}$. Therefore, for all $(\gamma,\delta)\in\Gamma_{f}$, $\alpha=\gamma(\alpha)=\delta(\sigma)\alpha$, so $\delta(\sigma)=1$, whence $\sigma_{|H}=\mathrm{id}$. Thus, $(\mathcal{B},\varepsilon_{f}):\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\mathrm{Hom}(G_{F},{\mathbb{F}}_{2})\times{\mathbb{C}}^{\times}$ is injective.
In particular, $\sigma\in\mathrm{Gal}(H/H_{0})\longmapsto\mathcal{B}(\sigma,\cdot)$ is injective, so that $|B|\geq[H:H_{0}]$, so we are done if $[H:H_{0}]>2$.
Now suppose that $[H:H_{0}]=2$ and let $\sigma_{0}$ be the unique non-trivial element of $\mathrm{Gal}(H/H_{0})$. For any $\sigma\in M$, the element $\alpha\in L^{\times}$ associated to $\sigma$ by Corollary 2 is such that $\frac{\alpha^{2}}{\varepsilon_{f}(\sigma)}\in F^{\times 2}$, so $\varepsilon_{f}(\sigma)\in L^{\times 2}$.
@@
$0\rightarrow\mathrm{Gal}(H/H_{0})\rightarrow\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\mathrm{Gal}(H_{0}/{\mathbb{Q}})\rightarrow 0$

### v3

@@ 2 Special elements in the image of the convolution of two modular forms
For (vi), let $r\geq 1$ be such that $\varepsilon_{f}^{2^{r}}$ is the trivial character. Let $(\gamma,\chi)$ be such that $f$ has an inner twist by $(\gamma,\chi)$. Then, by Lemma 1, $\chi$ is the product of a quadratic character and a power of $\varepsilon_{f}$, so $\chi^{2^{r}}$ is trivial, so $\mathrm{Gal}(H/{\mathbb{Q}})$ has exponent dividing $2^{r}$. Therefore, $d:=[H:{\mathbb{Q}}]$ is a power of $2$. In particular, $\varepsilon_{g}(G_{{\mathbb{Q}}})/\varepsilon_{g}(G_{H})$ and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/\tilde{\rho}_{g}(G_{H})$ are $2$-groups, so $\varepsilon_{g}(G_{H})$ contains an element of order $q$ and $q\mid|\tilde{\rho}_{g}(G_{H})|$.
We want to apply Proposition 2, but we need to rule out its exceptional cases. The first one cannot occur since, under our assumptions, $H:{\mathbb{Q}}]$ is not divisible by $3$. Now, assume that $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$: in particular, it has no non-trivial quotient which is a $2$-group. The image $I$ of $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a normal subgroup of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. Since $\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is onto, $d$ kills the quotient $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/I$, so $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/I$ is a $2$-group. Hence it is trivial, i.e. $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is onto: the second exception of Proposition 2 cannot occur and we are done. $\Box$
@@
Remark  When checking the Bloch–Kato conjecture, it seems more appropriate to use the $L$-function $L^{aut}(E,\rho,s)$ attached to the Rankin–Selberg convolution of the unitary automorphic representations attached to $f$ and $g$ (as in [2, Section 3.6]), since it admits a holomorphic continuation and satisfies a functional equation (with center $\frac{1}{2}$, see [6, Theorem 11.7.1]). We are using the imprimitive $L$-function in the statement above primarily for the sake of simplicity, since it is the one appearing throughout [10]. In fact, using the classical description of $L^{aut}(E,\rho,s)$ given in [12, Theorem 2.2], it is not difficult to check that the meromorphic continuation of $L(E,\rho,s)$ is indeed holomorphic near $s=1$ and that one has $L(E,\rho,1)\neq 0$ if and only if $L^{aut}(E,\rho,\frac{1}{2})\neq 0$.
@@ 3 Families of dihedral counter-examples: proof of Theorem Introduction
We first recall the following fact, already identified by Loeffler as an obstruction in [15, Proposition 4.1.1].
@@
Thus, we can (and do) assume that $M$ and $N$ both have distinct eigenvalues, respectively $m,m^{\prime}$ and $n,n^{\prime}$. Then $M\otimes N$ is diagonalizable with eigenvalues $mn,m^{\prime}n^{\prime},m^{\prime}n,mn^{\prime}$. Since $1=(mn)(m^{\prime}n^{\prime})=(mn^{\prime})(m^{\prime}n)$, there is never exactly one of these eigenvalues which is equal to one. $\Box$
As noted by Loeffler, this lemma implies that, for any couple $(f^{\prime},g^{\prime})$ of newforms, $T_{f^{\prime},g^{\prime},\mathfrak{p}}$ cannot satisfy (wE) if the product of the characters of $f^{\prime},g^{\prime}$ is trivial. That is why it is assumed in Question Introduction that $\varepsilon_{f}\varepsilon_{g}\neq 1$.
@@
    starts3-2special Let $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ be a prime ideal and $\sigma\in G_{{\mathbb{Q}}}$ be such that the kernel of $\sigma-1$ acting on $T_{f,g,\mathfrak{p}}$ is a line. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}})$ and $N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be the matrices of the action of $\sigma$ on $T_{f,\mathfrak{p}},T_{g}$. Then one has $\varepsilon(\sigma)=\varepsilon_{f}(\sigma)\varepsilon_{g}(\sigma)=\det{M}\det{N}\neq 1$ by Lemma 3, so $\sigma\notin G_{K}$. Since $g$ has real or complex multiplication by $K$, NN has trace zero, so, if $u\in{\mathbb{C}}^{\times}$ denotes a square root of $\det{N}$, $N$ is similar to $\Delta_{u,-u}$. This implies that $N^{2}$ is a scalar matrix. Moreover, $N$ is similar to N-N, so $M\otimes N$ is similar to $M\otimes(-N)=-(M\otimes N)$, so that $\ker(M\otimes N+1)$ is also a line.
@@
Proof. –  Suppose that $T_{f,g,\mathfrak{p}}$ satisfies (wE). Let $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ be such that $\ker(\sigma-1\mid T_{f,g,\mathfrak{p}})$ is a line. Let $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be associated to $\sigma$ by Corollary 2. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}}),N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be the matrices given by the action of $\sigma$ on $T_{f,\mathfrak{p}}$ and $T_{g}$, respectively. By point (ii) in the Remarks following Lemma 3, one has $\sigma\notin G_{K}$ and the matrix $N^{2}$ is scalar. In fact, one has by Cayley–Hamilton $N^{2}=(-\det{N})I_{2}=(-\varepsilon_{g}(\sigma))I_{2}=(-\varepsilon(\sigma)\varepsilon_{f}(\sigma)^{-1})I_{2}=\varepsilon_{f}(\sigma)^{-1}I_{2}$. Let $\omega\in L_{0}$ be a square root of $\varepsilon_{f}(\sigma)^{-1}$, then the two eigenvalues of $N$ are $\omega$ and $\omega$.
@@
By Corollary 2, we can write $M=\alpha^{-1}\varepsilon_{f}(\sigma)M^{\prime}$, where $M^{\prime}\in\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ has determinant $\frac{\alpha^{2}}{\varepsilon_{f}(\sigma)}$. Thus $\alpha\omega^{2}$ is an eigenvalue of $M^{\prime}\otimes\Delta_{\omega,-\omega}$, so that one of the eigenvalues of $M^{\prime}$ is $\pm\omega\alpha$. Since $(\pm\omega\alpha)^{2}=\det{M^{\prime}}$, one has $\pm\alpha\omega=\frac{1}{2}\operatorname{Tr}{M^{\prime}}\in F_{\mathfrak{p}}$, so $\mathcal{B}(\sigma,\mathrm{Frob}_{\mathfrak{p}})=1\neq\varepsilon(\sigma)$.
@@
Suppose conversely that $\mathcal{B}(\cdot,\mathrm{Frob}_{\mathfrak{p}})\neq\varepsilon$. Since $\mathcal{B}_{\mathfrak{p}}^{2}=1$ and $\varepsilon\neq 1$, $\varepsilon$ is not a power of $\mathcal{B}_{\mathfrak{p}}$, so $\ker{\mathcal{B}(\cdot,\mathrm{Frob}_{\mathfrak{p}})}$ is not contained in $\ker{\varepsilon}$. Hence, there exists $s\in G_{{\mathbb{Q}}}$ such that $\varepsilon(s)=-1$ and $\mathcal{B}_{\mathfrak{p}}(s)=1$. Let $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be attached to $s$ by Corollary 2 and $\nu:=\frac{\alpha^{2}}{\varepsilon_{f}(s)}\in\mathcal{O}_{F,\mathfrak{p}}^{\times}$. Since $\mathcal{B}_{\mathfrak{p}}(s)=1$, $\nu$ is the square of some $\nu^{\prime}\in\mathcal{O}_{F_{\mathfrak{p}}}^{\times}$. Therefore, $\varepsilon_{f}(s)^{-1}=\nu\alpha^{-2}$ is the square of $\omega:=\nu^{\prime}\alpha^{-1}\in L_{\mathfrak{p}}$.
@@
By Proposition 2 and Corollary 2, we can choose $\sigma_{1}\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\cap sG_{H}$ such that $\sigma_{1}$ acts on $T_{f,\mathfrak{p}}$ by the matrix $\alpha^{-1}\varepsilon_{f}(s)\begin{pmatrix}\nu^{\prime}&1\\ 0&\nu^{\prime}\end{pmatrix}$. Since $\varepsilon(\sigma_{1})=\varepsilon(s)=-1$, the matrix $\tilde{\rho}_{g}(\sigma_{1})\in\operatorname{PGL}_{2}(\mathcal{O}_{L_{0}})$ has order two. Since $\det{\rho_{g}(\sigma_{1})}=-\varepsilon_{f}(\sigma_{1})^{-1}=-\varepsilon_{f}(s)^{-1}$ and $p\neq 2$, $\rho_{g}(\sigma_{1})$ is $\operatorname{GL}_{2}(\mathcal{O}_{(L_{0}){\mathfrak{p}}})$-conjugate to $\Delta_{\omega,-\omega}$. Therefore, $\sigma_{1}$ acts on $T_{f,g,\mathfrak{p}}$ by $\begin{pmatrix}1&a&0&0\\ 0&1&0&0\\ 0&0&-1&\ast\\ 0&0&0&-1\end{pmatrix}$ with $a\in\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\times}$: hence $T_{f,g,\mathfrak{p}}/(\sigma_{1}-1)T_{f,g,\mathfrak{p}}$ is free of $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-rank one. $\Box$
@@
Proof. –  By Chebotarev and the previous Proposition, the answer to Question Introduction for $(f,g)$ is positive if and only if $\varepsilon$ cannot be written as a $\mathcal{B}(\cdot,\tau)$ for any $\tau\in G_{F}$. Because $\mathcal{B}$ is bilinear with values in a ${\mathbb{F}}_{2}$-vector space, the set $\{\mathcal{B}(\cdot,\tau)\mid\tau\in G_{F}\}$ is exactly the set of homomorphisms $f:\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\{\pm 1\}$ such that $f(M)=1$, and the conclusion follows. $\Box$
@@ 1. $(a)$
[…]
    weps-aspecial $k$ is even and $H$ is not totally real,
@@ 3. $(c)$
    weps-cspecial $k$ is odd, $H:H_{0}]=2$, and $\varepsilon_{f}$ does not take its values in $L^{\times 2}$ (this last condition is satisfied if $L$ does not contain a fourth root of unity $i$),
@@ 4. $(d)$
    weps-dspecial $k$ is odd, $H:H_{0}]=2$, $i\in L$, and for all $(\gamma,\delta)\in\Gamma_{f}$, $\gamma(i)=\delta(-1)i$,
@@ 5. $(e)$
Moreover, when $k$ is odd, $H:H_{0}]=2$, $i\in L$ but is not a value of $\varepsilon_{f}$, then the existence of such a $\varepsilon$ is equivalent to the last clause of $(e)$. Under the same assumptions, if $\varepsilon_{f}^{2}=1$, then $(d)$ and $(e)$ are equivalent.
Proof. –  Remember that $L$ is the number field generated by the coefficients of $f$, so it is a subfield of $\mathbb{C}$}. Note that $\mathcal{B}$ realizes an isomorphism from $\mathrm{Gal}(H/{\mathbb{Q}})/M$ to a subgroup $B$ of the ${\mathbb{F}}_{2}$-vector space $\mathrm{Hom}(G_{F},{\mathbb{F}}_{2})$, so $\mathrm{Gal}(H/{\mathbb{Q}})/M$ has exponent dividing $2$. Let $c\in\mathrm{Gal}(H/{\mathbb{Q}})$ be the image of the complex conjugation.
Let $\sigma\in M$ be such that $\varepsilon_{f}(\sigma)=1$: let $\alpha\in L^{\times}$ be associated to $\sigma$ by Corollary 2 and $\nu=\frac{\alpha^{2}}{\varepsilon_{f}(\sigma)}=\alpha^{2}\in F^{\times}$. Then, for all $\tau\in G_{F}$, $1=\mathcal{B}(\sigma,\tau)=\frac{\tau(\sqrt{\nu})}{\sqrt{\nu}}=\frac{\tau(\alpha)}{\alpha}$, hence $\alpha\in F^{\times}$. Therefore, for all $(\gamma,\delta)\in\Gamma_{f}$, $\alpha=\gamma(\alpha)=\delta(\sigma)\alpha$, so $\delta(\sigma)=1$, hence $\sigma\in G_{H}$. Thus, $(\mathcal{B},\varepsilon_{f}):\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\mathrm{Hom}(G_{F},{\mathbb{F}}_{2})\times{\mathbb{C}}^{\times}$ is injective.
In particular, $\sigma\in\mathrm{Gal}(H/H_{0})\longmapsto\mathcal{B}(\sigma,\cdot)$ is injective. In particular, if $k$ is even and $H$ is not totally real, then $c$ is non-trivial: since $\varepsilon_{f}(c)=1$, $c\notin M$, so there is a morphism $\mathrm{Gal}(H/{\mathbb{Q}})/M\rightarrow\{\pm 1\}$ sending $c$ to 1-1.
Suppose from now on that $k$ is odd. The goal is to show that there exists a non-trivial character $\mathrm{Gal}(H/{\mathbb{Q}})/\langle M,c\rangle\rightarrow\{\pm 1\}$. To do that, we need to show that $\mathrm{Gal}(H/{\mathbb{Q}})\supsetneq\langle M,c\rangle$, or in other words that $B\supsetneq\{1,\mathcal{B}(c,\cdot)\}$. Since $\sigma\in\mathrm{Gal}(H/H_{0})\longmapsto\mathcal{B}(\sigma,\cdot)$ is injective, one has $|B|\geq[H:H_{0}]$, so we are done if $H:H_{0}]>2$.
Suppose from now on that $H:H_{0}]=2$ and let $\sigma_{0}$ be the unique non-trivial element of $\mathrm{Gal}(H/H_{0})$. For any $\sigma\in\mathrm{Gal}(H/{\mathbb{Q}})$, $\sigma\in M$ if and only if the element $\alpha\in L^{\times}$ associated to $\sigma$ by Corollary 2 is such that $\frac{\alpha^{2}}{\varepsilon_{f}(\sigma)}\in F^{\times 2}$, which is the case if and only if $\sqrt{\varepsilon_{f}(\sigma)}$ lies in $L^{\times}$ and is attached to $\sigma$ by Corollary 2.
@@
$0\rightarrow\mathrm{Gal}(H/H_{0})\rightarrow\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\mathrm{Gal}(H_{0}/{\mathbb{Q}})\rightarrow 0.$